OpenStudy (anonymous):

@deana99 @TheRaggedyDoctor

OpenStudy (theraggedydoctor):

Hayy

OpenStudy (anonymous):

hey lol

OpenStudy (anonymous):

So which one is it

OpenStudy (theraggedydoctor):

I say C, she says A. Personally, I'm not 100% positive on this one, but I'm preeeeetty sure.

OpenStudy (theraggedydoctor):

I dunno tho

OpenStudy (anonymous):

A music competition on television had five elimination rounds. After each elimination, only half of the contestants were sent to the next round. The table below shows the number of contestants in each round of the competition: x 1 2 3 4 5 f(x) 64 32 16 8 4 Compute the average rate of change of f(x) from x = 3 to x = 5 and describe what it represents. -6 contestants per round, and it represents the number of contestants who will reach the final round -28 contestants per round, and it represents the number of contestants who will reach the final round -6 contestants per round, and it represents the average rate at which the number of contestants decreased from the third round to the fifth round -28 contestants per round, and it represents the average rate at which the number of contestants decreased from the third round to the fifth round

OpenStudy (anonymous):

yeah :) hold on

OpenStudy (deana99):

well I'm wrong so Im going to shut up and leave it to him

OpenStudy (anonymous):

The number of subscribers y to a website after t years is shown by the equation below: y = 50(1.75)t Which conclusion is correct about the number of subscribers to the website? It increased by 75% every year. It decreased by 75% every year. It increased by 50% every year. It decreased by 50% every year.

OpenStudy (theraggedydoctor):

How many more questions? And this one's A.

OpenStudy (anonymous):

idk let me check this is a practice thing to tell us what is going to be on the test and it is extra credit so if i get everything right i get extra credit

OpenStudy (anonymous):

which is really nice because i need extra credit and we have about 15 more to go srry than you so much for your cooperation ;)

OpenStudy (anonymous):

The table below shows the functions to represent the sale f(t), in thousands, of four companies in t years: Company A B C D Sale f(t) = 1.05t f(t) = 0.75t f(t) = 1.212t f(t) = 0.9712t Which of the following companies showed an exponential growth in sale? Company A and Company C Company B and Company D Company C and Company D Company A and Company B

OpenStudy (anonymous):

company b and d right @TheRaggedyDoctor @deana99

OpenStudy (theraggedydoctor):

Yup, that's right

OpenStudy (theraggedydoctor):

I'm not sure if I'll be able to do all 15, btw. Lol

OpenStudy (anonymous):

Henry is taking a medicine for a common cold. His body metabolizes the medicine over a certain time interval. The table below shows the amount of medicine f(t), in mg, that was present in Henry's body after time t: t (hours) 1 2 3 4 5 f(t) (mg) 282 265.08 249.18 234.22 220.17 Greg was administered 200 mg of the same medicine. The amount of medicine in his body f(t) after time t is shown by the equation below: f(t) = 200(0.88)t Which statement best describes the rate at which Henry’s and Greg's bodies eliminated the medicine? Henry's body eliminated the antibiotic faster than Greg's body. Henry's body eliminated the antibiotic at the same rate as Greg's body. Henry's body eliminated the antibiotic at half of the rate at which Greg's body eliminated the antibiotic. Henry's body eliminated the antibiotic at one-fourth of the rate at which Greg's body eliminated the antibiotic.

OpenStudy (anonymous):

lol ok at least try :D srry for being so forceful I just want to get it done

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

A?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The price of gold has increased by 35% per year from 2000. In the year 2000, Harry bought a gold ring for $590. Which of the following functions f(x) can be used to represent the price of the ring x years after 2000? f(x) = 590(1.35)x f(x) = 590(0.65)x f(x) = 35(0.41)x f(x) = 35(1.59)x

OpenStudy (theraggedydoctor):

A

OpenStudy (anonymous):

The average annual salary of the employees of a company in the year 2005 was $80,000. It increased by the same factor each year and in 2006, the average annual salary was $88,000. Let f(x) represent the average annual salary, in thousand dollars, after x years since 2005. Which of the following best represents the relationship between x and f(x)? f(x) = 88(0.88)x f(x) = 88(1.1)x f(x) = 80(0.88)x f(x) = 80(1.1)x

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

D.

OpenStudy (anonymous):

Which statement best describes the effect of replacing the function f(x) = 2x + 2 with the function g(x) = 2x - 3? The graph shifts 1 unit left. The graph shifts 5 units left. The graph shifts 5 units right. The graph shifts 1 unit right.

OpenStudy (anonymous):

@TheRaggedyDoctor

OpenStudy (theraggedydoctor):

5 units right, sorry.

OpenStudy (deana99):

can you take my name off here please

OpenStudy (anonymous):

??? @deana99

OpenStudy (theraggedydoctor):

Or make a new question! I wanna be green. All my friends are green. Meh.